OpenStudy (dancergirl45):

NEED HELP FAST WILL FAN AND MEDAL Solve for x: 5|2x − 2| + 8 = 18

OpenStudy (study312):

I looked up your question and found this if it makes any sense 5 |2x-2|+8=18 5|2x-2|= 18-8 5|2x-2|=10 5/5 |2x-2|=10/5 |2x-2|=2 2|x-1|=2 |x-1|=1 x-1= 1 or x-1 = -1 x=2 0r x=0 check: If x=2, LHS:5|4-2|+8=10+8=18=RHS If x=0, LHS= 5|0-2|+8=10+8=18=RHS Ans: x=2, x=0

jhonyy9 (jhonyy9):

nice work @Study312